OpenStudy (anonymous):

H(x) = e^(3x-2) If h(x)=f(g(x)), find f(x) and g(x) using the Chain Rule?

OpenStudy (anonymous):

g(x)=3x-2 and f(x)=e^g(x)

OpenStudy (anonymous):

g(x) = 3x-2 f(x) = e^x

OpenStudy (anonymous):

w=g(x)=3x-2 v=f(x)=e^w h(x)=f(g(x)) h(x)=e^w h(x)=e^(3x-2)